# Domain decomposition of the modified Born series approach for large-scale wave propagation simulations

### v2

@@ Abstract
The modified Born series method is a fast and accurate method for simulating wave propagation in complex structures. Currently, its main limitation is that the simulation size is limited by the working memory of a single computer or graphics processing unit (GPU). Here, we present a domain decomposition method that removes this limitation. We show how to decompose large problems over subdomains while maintaining the method’s accuracy, memory efficiency, and guaranteed monotonic convergence. We demonstrate our approach by solving the Helmholtz problem for a complex structure of size $315\times 315\times 315$ wavelengths in just $1.4$ hours on a dual-GPU system.
@@ 1 Introduction
[…]
Unlike widely used methods such as the finite-difference time-domain (FDTD) [4, 5, 6, 7] and pseudospectral time-domain (PSTD) [8, 9, 10] methods, the MBS does not rely on finite difference approximations. This is an important benefit since the finite difference approximations accumulate as the wave propagates, easily causing errors of more than 1% per propagated wavelength [11], meaning that the relative phase of two beams can be off by $\pi$ already after 50 wavelengths. The numerical dispersion of finite approximations is well-studied in literature [8, 12, 13, 14].
@@
An additional benefit of MBS is that it requires far fewer iteration steps than FDTD or PSTD methods [2]. These combined benefits currently make the MBS one of the most efficient methods to solve Helmholtz and Maxwell’s equations [2, 15], with subsequent extensions to birefringent media [16], the inverse scattering problem [17], and applications to photoacoustics [18] and geophysics [19]. Transient wave phenomena, which are especially important in sound and seismic waves, can be simulated by solving for each frequency component individually, as is common in frequency-domain methods [20, 21, 22, 23, 24, 25]. This approach has the advantage of being able to accurately describe physical dispersion effects (i.e. wavelength-dependent propagation speed).
@@
Osnabrugge et al. [3] showed that the MBS could be executed on a graphics processing unit (GPU) efficiently, solving Maxwell’s equations in a structure of $28\times 28\times 28$ wavelengths in just 3 seconds, representing a factor of $50$ speed-up compared to performing the simulations on a central processing unit (CPU). Recently, Valantinas and Vettenburg [26] solved the Helmholtz equation for a system of $2.1\cdot 10^{7}$ cubic wavelengths by mapping the steps of the MBS to layers in an artificial neural network framework, leveraging the extensive software and hardware acceleration available for these applications. However, GPUs have a limited amount of expensive working memory. They cannot be expanded as flexibly as the main working memory, adding a severe limitation to the simulation size [27, 28].
For FDTD, this problem has been solved by domain decomposition, i.e., by distributing the computations over multiple PCs or GPUs, each solving a subdomain of the original problem [29, 30, 31, 32]. This approach enables simulations of arbitrary size, limited only by the total memory available in a cluster or multi-GPU PC.
Here, we present a domain decomposition approach for the MBS to solve the Helmholtz equation. Our approach, based on a non-overlapping domain decomposition strategy [33, 34, 35, 36], maintains the key benefits of MBS: low memory use, high accuracy, and guaranteed convergence, while minimising the overhead on memory use and computation time. Whereas it is common to explicitly formulate boundary conditions for the wave equation [36, 37, 38], our approach implicitly handles both the external boundary conditions and the boundary conditions at the domain boundaries, allowing both periodic and non-periodic boundaries. At the exterior, these boundaries are combined with a thin absorbing anti-reflection layer [3], eliminating the need for computationally expensive padding with thick absorbing boundary layers [8, 37].
@@
We demonstrate the domain decomposition approach on a dual-GPU system through a large 3D simulation of $3.1\cdot 10^{7}$ cubic wavelengths, a size that is 1.93×1.93$\times$1.93 × larger than the maximum that could be accommodated on a single GPU and took only $1.4$ hours to complete. An open-source implementation in Python is available on GitHub [39].
@@
We start in Section 2 by summarising the MBS approach. In Section 3, we introduce the proposed domain decomposition method. We then demonstrate the approach and analyse its efficiency and accuracy (Section 4). Concluding remarks are provided in Section 5.
@@ 2 Modified Born series
$\left[\nabla^{2}+{k}^{2}(\boldsymbol{r})\right]{x}(\boldsymbol{r})=-{y}(\boldsymbol{r}),$ (1)
where ${k}(\boldsymbol{r})$ is the wavenumber, ${x}(\boldsymbol{r})$ is the field, and ${y}(\boldsymbol{r})$ is the source term as a function of the position vector $\boldsymbol{r}$. For convenience, we introduce the notation
@@
[…]
The key step in both the modified [2] and original [41] Born series formalisms is to split $A=L+V$, with $L\coloneqq c\left[\nabla^{2}+k_{0}^{2}\right]$ corresponding to wave propagation in a homogeneous medium with background wavenumber $k_{0}$, and $V=c\left[{k}^{2}(\boldsymbol{r})-k_{0}^{2}\right]$ the scattering potential.
@@
In contrast to the original Born series, which diverges for large or strongly scattering media [45], the MBS converges monotonically to the solution of Eq. (1) provided that $\alpha\leq 1$ and $ic$ is a real number less than $1/|k^{2}(\boldsymbol{r})-k_{0}^{2}|_{\infty}$ [2], where $|\cdot|_{\infty}$ denotes the maximum value over all $\boldsymbol{r}$. We can truncate this iteration when the norm of the residual $\Gamma^{-1}\left(\boldsymbol{y}-A\boldsymbol{x}^{(n)}\right)$ drops below a certain threshold.
@@
$c=-\frac{0.95i}{\left\lVert{k}^{2}(\boldsymbol{r})-k_{0}^{2}\right\rVert_{\infty}},\qquad\alpha=0.75.$ (4)
The splitting $A=L+V$ is visualised in Fig. 1a-c, where the operators are represented as matrices for illustrative purposes. In practice, these operators are not implemented as matrices since the full matrices would be too large to store in memory. The alternative of using sparse matrices would require replacing $L$ by a matrix with only a few diagonals, which is precisely the finite difference approximation that causes numerical dispersion and thus should be avoided. Instead, $L$ can be implemented as a convolution with the angular spectrum kernel [46, 47]
$L=c{\mathcal{F}}^{-1}\left(-\left\lVert\boldsymbol{p}\right\rVert^{2}+k_{0}^{2}\right)\mathcal{F},$ (5)
where $\mathcal{F}$ and ${\mathcal{F}}^{-1}$ denote the forward and inverse Fourier transform, respectively, and $\boldsymbol{p}$ denotes the Fourier-space coordinate vector. In practice, a fast Fourier transform is used to evaluate this convolution efficiently. As shown in Fig. 1d, this approach does introduce wrapping artefacts. These artefacts can be significantly reduced by using an alternating sequence of offset Fourier transforms [3] and eliminated using our proposed method (see Section 3). Either of these methods is then combined with a thin absorbing boundary layer to ensure convergence [3].
Similar to Eq. (5), we can implement $(L+I)^{-1}$ as a (fast) convolution. Substituting $\Gamma^{-1}=(I-V)(L+I)^{-1}$ and $A=(L+I)-(I-V)$ into Eq. (3) gives the iteration [2]
$\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}-\alpha(I-V)\left(\boldsymbol{x}^{(n)}-(L+I)^{-1}\left[\boldsymbol{y}+(1-V)\boldsymbol{x}^{(n)}\right]\right),$ (6)
@@
Figure 1: Matrix representation of splitting operator $A$ into $L+V$ for a 1D system. (a) Operator $A$ is split into (b) $L$ containing the Laplacian operator, a band along the diagonal, and (c) $V$, a diagonal matrix containing the scattering potential. (d) The operator $L$ with wraparound artefacts resulting from implementing $L$ as a fast convolution (Eq. (5)).
@@ 3 Domain decomposition of the modified Born series
We start by decomposing the linear system $A\boldsymbol{x}=\boldsymbol{y}$ over multiple subdomains. For illustrative purposes, we decompose a 1D problem into two subdomains of equal size $N$. The same strategy extends to 2D and 3D problems, with any number of subdomains along any dimension and domains of unequal size. The operator $A$ for this 1D problem can be represented as a block matrix:
@@
where the blocks $A_{11},A_{12},A_{21}$, and $A_{22}$ are of size ${N\times N}$, and the vectors $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\boldsymbol{y}_{1}$, and $\boldsymbol{y}_{2}$ are of length $N$. Here, $A_{11}$ and $A_{22}$ are matrices operating on the two subdomains, while the blocks $A_{12}$ and $A_{21}$ represent information transfer between the subdomains. The matrix representation of the decomposition of $A$ over two subdomains is shown in Fig. 2a.
@@
Figure 2: Matrix representation of domain decomposition. (a) $A$ (same as Fig. 1a) is decomposed over two subdomains, shown by the grey lines. (b) Decomposition of $L$, implemented as a fast convolution (Eq. (5)) over each subdomain, thus containing wraparound artefacts. The off-diagonal communication blocks are empty. (c) $V\coloneqq A-L$ contains the scattering potential on the diagonal and the wraparound and communications corrections.
@@
An essential generalisation of the MBS introduced in [40] is that we have complete freedom to choose the splitting $A=L+V$, as long as the scaling factor $c$ can be selected such that $\lVert V\rVert<1$. We use this freedom to implement our domain decomposition framework.
@@
This freedom enables us to implement $L$ as a fast convolution over each subdomain (see Fig. 2b). Representing $L$ and $V$ in block matrix form
$L=\begin{bmatrix}L_{11}&0\\ 0&L_{22}\end{bmatrix},\quad V=\begin{bmatrix}A_{11}-L_{11}&A_{12}\\ A_{21}&A_{22}-L_{22}\end{bmatrix},$ (8)
we see that $L$ is a domain-local operator and $V$ is mostly domain-local, with small, explicit contributions for transfer between domains (represented by $A_{12}$ and $A_{21}$, also seen in Fig. 2c). It is important to note that the system being solved here is still the same as in Eq. (2).
In $L$ (Eq. (8)), the diagonal blocks $L_{11}$ and $L_{22}$ (visualized in Fig. 2b) contain wraparound artefacts, similar to Fig. 1d. These wraparound artefacts appear in $V$ with the opposite sign, as $V\coloneqq A-L$. $V$ now has three contributions (Fig. 2c). The domain-local part constitutes the original scattering potential on the diagonal and the correction for the wrapping artefacts caused by the fast convolution, which we call $C_{11}$ and $C_{22}$. The non-local part of $V$ is the off-diagonal blocks $A_{12}$ and $A_{21}$ representing communication between the subdomains.
The interior subdomain communication and exterior boundary corrections are based on these matrices $C_{11}$, $C_{22}$, $A_{12}$, and $A_{21}$ combined with a thin absorbing boundary layer [3] on the exterior. For the exterior boundary, the matrices $C_{11}$ and $C_{22}$ are used to tackle the wrapping artefacts, which allows for only a thin absorbing boundary layer to minimize reflection artefacts. For the interior boundaries, we use the matrices $C_{11}$ and $C_{22}$ to tackle the wrapping artefacts at the interface between subdomains, and the matrices $A_{12}$ and $A_{21}$ for communication between the subdomains.
Thus, our approach does not require explicit steps to guarantee continuity at the subdomain boundaries but instead results in the naturally occurring boundary conditions at the interface through the matrices $A_{12}$ and $A_{21}$. The matrices $C_{11}$, $C_{22}$, $A_{12}$, and $A_{21}$ can all be constructed directly by comparing two situations: 1) the angular spectrum kernel applied to a point source using a linear convolution (Fig. 3a), and 2) the same kernel applied using a fast convolution on a subdomain (Fig. 3b). The difference between these two curves gives the required wrapping correction (which corresponds to a column of matrix $C_{11}$) and the transfer to the second subdomain (a column of matrix $A_{12}$). By repeating this process for different positions of the point source and thus different columns of the matrices, we recover matrices $C_{11}$ and $A_{12}$. Due to the symmetry of the angular spectrum kernel, the bottom-left part of matrix $C_{11}$ is equal to $-A_{12}$. The top-right part of matrix $C_{11}$ is given by the transpose $-A_{12}^{T}$.
@@
Figure 3: Generation of the matrix $A_{12}$ for wrapping and transfer corrections. (a) Field after convolution of the Laplacian with a point source. (b) Field when the fast convolution is computed only over Subdomain 1 (white background). The difference between the fields in (a) and (b) is shown by the dashed red lines and gives columns of (c) the (truncated) correction matrix $A_{12}$.
Figure 3c visualizes the bottom-left part of the domain-transfer matrix $A_{12}$ (also see Fig. 2c), which shows that the magnitude of the elements decreases rapidly away from the subdomain border. Therefore, we can truncate matrices $C$ and off-diagonal blocks of $A$ to only consider $t\ll N$ edge points near the boundaries. This way, the overhead of these additional computations is minimised.
@@
Figure 4: Using truncated corrections in $V$ to reconstruct $A$. (a) Matrix representation of $A$ decomposed over two subdomains, demarcated by the dashed grey lines, but with truncated corrections. (b) $L$ decomposed over the subdomains remains unchanged and is the same as in Fig.2b. (c) $V$ with truncated corrections for wrapping and transfer, visible as square blocks of side $t=6$ in the corners of subdomain blocks. Note that the colour scale is logarithmic in both the positive and negative directions from 0.
@@
Fig. 4 visualises the truncated corrections in $V$ for the same example as Fig. 2. The operator $A$ obtained using the truncated corrections is shown in Fig. 4a. The decomposed operator $L$ (Eq. (8)) shown in Fig. 4b remains unchanged irrespective of the truncation and is the same as in Fig. 2b. Fig. 4c shows the matrix representation of the operator $V$ obtained using the truncated corrections instead of the full corrections as in Fig. 2d. The effect of the truncation is visible in Fig. 4c as square blocks of size $t=6$.

### v3

@@ Abstract
The modified Born series (MBS) is a fast and accurate method for simulating wave propagation in complex structures. In the current implementation of the MBS, the simulation size is limited by the working memory of a single computer or graphics processing unit (GPU). Here, we present a domain decomposition method that enhances the scalability of the MBS by distributing the computations over multiple GPUs, while maintaining its accuracy, memory efficiency, and guaranteed monotonic convergence. With this new method, the computations can be performed in parallel, and a larger simulation size is possible as it is no longer limited to the memory size of a single computer or GPU. We show how to decompose large problems over subdomains and demonstrate our approach by solving the Helmholtz problem for a complex structure of $3.28\cdot 10^{7}$ cubic wavelengths ($320\times 320\times 320$ wavelengths) in just $45$ minutes with a dual-GPU simulation.
@@ 1 Introduction
Wave propagation simulations have many applications, ranging from nanophotonics to geophysics. Unfortunately, computing accurate solutions to wave equations in large heterogeneous media is highly time-consuming. A great number of numerical methods are available for solving wave equations in heterogeneous media [1]. The modified Born Series (MBS) [2, 3] is one of the fastest and most accurate methods [4, 5].
@@
The MBS is a frequency domain method, where the structure is represented on a regular grid. This grid-based approach allows arbitrarily complex structures to be defined and enables an efficient computation of the iteration steps. Transient wave phenomena, which are especially important in sound and seismic waves, can be simulated by solving for each frequency component individually, as is common in frequency-domain methods [25, 26, 27, 28, 29, 30]. This approach has the advantage of being able to accurately describe physical dispersion effects (i.e., wavelength-dependent propagation speed).
A key benefit of the MBS is that it does not rely on finite difference approximations. This sets it apart from widely used grid-based methods such as the finite-difference time-domain (FDTD) [6, 7, 8, 9] and pseudospectral time-domain (PSTD) [10, 11, 12] that rely on finite difference approximations for the spatial (FDTD) and temporal (FDTD and PSTD) derivatives. These approximations introduce errors that accumulate as the wave propagates [4, 10, 14, 15, 16]. Although it is possible to reduce these errors by decreasing the voxel size (for FDTD) and the time step (for FDTD and PSTD) [2, 17, 18, 19, 20], this comes at a correspondingly steep cost in memory usage and computation time while giving a limited benefit; for example, even at a sampling of 10 points per wavelength, the phase velocity error of FDTD is a few percent [13], resulting in a phase error of over $\pi$ after propagating roughly a dozen of wavelengths. Similarly, accumulating errors in the time derivative limit PSTD to simulations far less than 100 wavelengths, even at a time step of 1/10 per propagated wavelength [2].
MBS does not suffer from these limitations. Like PSTD, it only requires the spatial sampling to be fine enough to sample the spatial frequencies in the simulated field, i.e., the Nyquist sampling limit for the simulation, allowing large simulation domains to fit in memory. In addition, since the time derivative is also not approximated, MBS requires far fewer (and approximately equally fast) iterations than PSTD. Where PSTD needs 10 or more steps per wavelength, MBS covers the simulation domain in steps of multiple wavelengths at once, a phenomenon known as pseudo-propagation [2].
These combined benefits currently make the MBS one of the most efficient methods to solve Helmholtz and Maxwell’s equations [2, 4, 5], with subsequent extensions to birefringent media [21], the inverse scattering problem [5], and applications to sound and seismic waves [22, 23, 24]. In experiments, the MBS has been demonstrated to be more than three orders of magnitude faster than FDTD [4] and PSTD [2].
@@
A significant speed increase can be achieved by executing the simulations on a graphics processing unit (GPU). For example, Osnabrugge et al. [3] used MBS on a GPU to solve Maxwell’s equations in a structure of $19.7\cdot 10^{3}$ cubic wavelengths (sampled at $5$ points per wavelength) in just 3 seconds, representing a factor of $50$ speed-up compared to performing the simulations on a central processing unit (CPU) (in both cases, the Mie sphere simulations were run until the maximum accuracy, a relative error of $0.014$ with the analytical solution, was reached). Recently, Valantinas and Vettenburg [31] solved the Helmholtz equation for a system of $2.1\cdot 10^{7}$ cubic wavelengths sampled at $3$ points per wavelength, reaching a residue of $10^{-4}$ in $13$ minutes and $4270$ iterations, by mapping the steps of the MBS to layers in an artificial neural network framework, leveraging the memory efficiency of the MBS and the extensive software and hardware acceleration available for these applications. For comparison, one of the largest recent FDTD simulations was a much smaller $4.6\cdot 10^{5}$ cubic wavelengths (sampled at $18$ points per wavelength), which was run for $6235$ time steps on an elastic network of an unspecified number of NVIDIA GPUs and took $5$ minutes [32]. However, while GPUs offer significantly faster performances than CPUs, they have a limited amount of working memory. They cannot be expanded as flexibly as the main working memory, adding a severe limitation to the simulation size [33, 34]. For FDTD, this problem of limited memory has been solved by domain decomposition, i.e., by distributing the computations of a huge problem over multiple PCs or GPUs, each solving a subdomain of the original problem [35, 36, 37, 38, 39, 40].
Here, we present a method to parallelise the MBS by distributing the computations over multiple GPUs. Our non-overlapping domain decomposition strategy [35, 36, 37, 38, 39] is derived directly from a block operator decomposition of the Helmholtz equation. This eliminates the need for the commonly used explicit subdomain boundary conditions (e.g., [38, 41, 40]), and makes it possible to maintain the key benefits of the MBS: low memory use, high accuracy, and guaranteed monotonic convergence, while minimising the overhead on memory use and computation time. We show that the rate of convergence does not deteriorate when the number of subdomains grows, making our method scalable according to the definition in Refs. [38, 39].
@@
We demonstrate the domain decomposition approach on a GPU system through a large 3D simulation of $3.27\cdot 10^{7}$ cubic wavelengths sampled at $4$ points per wavelength, a size that is $1.95\times$ larger than the maximum that could be accommodated on a single GPU, and takes only $45$ minutes to reach a relative residual of $10^{-6}$. With this new framework, a larger simulation size is possible as it is no longer limited to the memory size of a single graphics card. Further, the computations can be parallelised and scaled to even larger domains by adding more GPUs. An open-source implementation in Python is available on GitHub [42].
@@
We start in Section 2 by summarising the MBS approach. In Section 3, we introduce the proposed domain decomposition method and discuss its implementation, computational complexity and memory usage. We then demonstrate the approach and analyse its efficiency and accuracy (Section 4). Concluding remarks are provided in Section 5.
@@ 2 Modified Born series
$\left(\nabla^{2}+{k}^{2}\right)\psi=-S,$ (1)
with the Laplacian $\nabla^{2}$, spatially varying wavenumber k{k}, field $\psi$, and source $S$. For convenience, we introduce the operator notation
@@
where $A\coloneqq c\left(\nabla^{2}+k^{2}\right)$ is a linear operator, $\boldsymbol{x}\coloneqq\psi$ and $\boldsymbol{y}\coloneqq-cS$ are the field and scaled source sampled on a regular grid, and $c$ is a scaling factor discussed below.
The key step in both the modified [2] and original [44] Born series formalisms is to split $A$ into $L+V$, with $L\coloneqq c\left[\nabla^{2}+k_{0}^{2}\right]$ corresponding to wave propagation in a homogeneous medium with background wavenumber $k_{0}$, and $V=c\left[{k}^{2}-k_{0}^{2}\right]$ the scattering potential.
@@
In contrast to the original Born series, which diverges for large or strongly scattering media [48], the MBS converges monotonically to the solution of Eq. (1) provided that $\alpha\leq 1$ and $ic$ is a real number less than $1/\left\lVert k^{2}-k_{0}^{2}\right\rVert_{\infty}$ [2], where $\left\lVert\cdot\right\rVert_{\infty}$ denotes the maximum value over all $\boldsymbol{r}$. We truncate this iteration when the norm of the residual $\Gamma^{-1}\left(\boldsymbol{y}-A\boldsymbol{x}^{(n)}\right)$ drops below a certain threshold.
Figure 1: Convergence of the MBS for a problem of size $50\times 50\times 50$ wavelengths sampled at $4$ points per wavelength (details in Sections 4 and 4.1) for different refractive index values of the scattering medium (the background refractive index is kept 11). The real part ranges from $1$ to 55, and the imaginary part (extinction coefficient) from 0 to $9$. The (a) number of iterations and (b) simulation time to converge to a relative residual of $10^{-6}$ increase with scattering contrast as expected.
@@
$c=-\frac{0.95i}{\left\lVert{k}^{2}-k_{0}^{2}\right\rVert_{\infty}},\qquad\alpha=0.75,$ (4)
which shows that $c$ is inversely proportional to the scattering contrast in the simulation. The number of iterations of the MBS is, in turn, inversely proportional to the scaling factor $c$, so the number of iterations increases with the scattering contrast [2]. Figure 1 demonstrates this effect for a simulation of scalar wave propagation in a collection of spheres (details in Sections 4 and 4.1). The refractive index (or relative wave velocity) of the scattering medium ranges from $1$ to $5$ for the real part and from 0 to $9$ for the imaginary part, while the background refractive index is $1$. It can be seen from the figure that the MBS converges for all refractive index values considered, and the number of iterations (and simulation time) increases with the scattering contrast.
Figure 2: Matrix representation of splitting operator AA into $L+V$ for a 1D system. (a) Operator AA is split into (b) $L$ containing the Laplacian operator, a band along the diagonal, and (c) $V$, a diagonal matrix containing the scattering potential. (d) The operator $L$ with wraparound artefacts resulting from implementing $L$ as a fast convolution (Eq. (5)).
The splitting $A=L+V$ is visualised for a 1D problem in Fig. 2a-c, where the operators are represented as matrices for illustrative purposes. In practice, these operators are not implemented as matrices, as the full matrices would be too large to store in memory. The alternative of using sparse matrices would require replacing $L$ by a matrix with only a few diagonals, which is precisely the finite difference approximation that causes numerical dispersion and thus should be avoided. Instead, $L$ can be implemented as a convolution with the angular spectrum kernel [49, 50]
$L=c{\mathcal{F}}^{-1}\left(-\left\lVert\boldsymbol{p}\right\rVert_{2}^{2}+k_{0}^{2}\right)\mathcal{F},$ (5)
where $\mathcal{F}$ and ${\mathcal{F}}^{-1}$ denote the forward and inverse Fourier transform, respectively, $\boldsymbol{p}$ denotes the Fourier-space coordinate vector, and $\left\lVert\cdot\right\rVert_{2}$ is the Euclidean norm. In practice, a fast Fourier transform (FFT) is used to efficiently evaluate this convolution. As shown in Fig. 2d, this approach does introduce wrapping artefacts. These artefacts can be significantly reduced by using an alternating sequence of offset Fourier transforms [3] and eliminated using our proposed method (see Section 3). Either of these methods can then be combined with a thin anti-reflection layer [3] to prevent reflections at the domain boundaries.
Similar to Eq. (5), we can implement $(L+I)^{-1}$ as a (fast) convolution. Substituting $\Gamma^{-1}=(I-V)(L+I)^{-1}$ and A=(L+I)(IV)A=(L+I)-(I-V) into Eq. (3) gives the iteration [2]
$\boldsymbol{x}^{(n+1)}=\boldsymbol{x}^{(n)}-\alpha(I-V)\left(\boldsymbol{x}^{(n)}-(L+I)^{-1}\left[\boldsymbol{y}+({I}-V)\boldsymbol{x}^{(n)}\right]\right),$ (6)
@@ 3 Domain decomposition of the modified Born series
To parallelise the MBS so that it can be executed on multiple GPUs simultaneously, we start by decomposing the linear system $A\boldsymbol{x}=\boldsymbol{y}$ (Eq. (2)) over multiple subdomains. For illustrative purposes, we decompose a 1D problem into two subdomains of equal size $N$. The same strategy extends to 2D and 3D problems, with any number of subdomains along any dimension and domains of unequal size. The operator $A$ for this 1D problem can be represented as a block decomposition:
@@
where the blocks $A_{11},A_{12},A_{21}$, and $A_{22}$ are of size $N\times N$}, and the vectors $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\boldsymbol{y}_{1}$, and $\boldsymbol{y}_{2}$ are of length $N$. Here, the blocks $A_{11}$ and $A_{22}$ operate on subdomains 1 and 2, respectively, while the blocks $A_{12}$ and $A_{21}$ represent communication between the subdomains. The decomposition of $A$ over two subdomains is visualised in Fig. 3a. Note that, as before, these blocks are linear operators and not stored as matrices explicitly.
@@
Figure 3: Matrix representation of the block decomposition of Eq. (2). (a) AA (same as Fig. 2a) is decomposed over two subdomains, shown by the dashed grey lines. (b) Decomposition of $L$, implemented as a fast convolution (Eq. (5)) over each subdomain, thus containing wraparound artefacts. The off-diagonal (communication) blocks are empty. (c) $V\coloneqq A-L$ has diagonal blocks containing the scattering potential on the diagonal elements and wrapping artefacts from $L$ but with the opposite sign, and off-diagonal blocks representing communication between the subdomains.
@@
An essential generalisation of the MBS introduced in [43] is that we have complete freedom to choose the splitting $A=L+V$, as long as the scaling factor $c$ can be selected such that $\left\lVert V\right\rVert<1$, where $\left\lVert\cdot\right\rVert$ denotes the operator norm, i.e., its largest singular value. We use this freedom to implement our domain decomposition framework.
@@
This freedom enables us to implement $L$ as a fast convolution over each subdomain (see Fig. 3b), with $V\coloneqq A-L$ [43]. Representing $L$ and $V$ as a block decomposition
$L=\begin{bmatrix}L_{11}&0\\ 0&L_{22}\end{bmatrix},\quad V=\begin{bmatrix}A_{11}-L_{11}&A_{12}\\ A_{21}&A_{22}-L_{22}\end{bmatrix}\coloneqq\begin{bmatrix}V_{11}&A_{12}\\ A_{21}&V_{22}\end{bmatrix}$, (8)
we see that $L$ (visualised in Fig. 3b) is a domain-local operator, with the diagonal blocks $L_{11}$ and $L_{22}$ operating on the subdomains, and contains wraparound artefacts similar to Fig. 2d, but now over the subdomains. $V$ now has both diagonal and off-diagonal blocks (visualised in Fig. 3c). The diagonal blocks $V_{11}$ and $V_{22}$ contain, as before, the scattering potential on the diagonal elements (as in Fig. 2c), and now also the wrapping artefacts from $L_{11}$ and $L_{22}$ but with the opposite sign. The off-diagonal blocks $A_{12}$ and $A_{21}$ represent communication between the subdomains.
Unlike common domain decomposition methods [35, 36, 38, 40, 51, 52, 53], our approach does not require the explicit specification of boundary conditions at the subdomain boundaries, because when added up, $L+V$ is still exactly identical to the original system (Eq. (2)).
@@
Figure 4: Generation of the off-diagonal blocks $A_{12}$ and $A_{21}$ using the angular spectrum kernel. $A_{21}=A_{12}^{T}$, so we only show $A_{12}$. (a) Field after convolution of the Laplacian with a point source. (b) Field when the fast convolution is computed only over Subdomain 1 (white background). The difference between the fields in (a) and (b) is shown by the dashed red lines and gives columns of (c) the (truncated) block $A_{12}$.
In our framework, we do need a way to efficiently compute the off-diagonal elements in $V$. We realise that these elements represent the difference between the kernel applied to a point source in an infinite domain (Fig. 4a) versus a periodic subdomain (Fig. 4b). Fig. 4c visualises this process repeated for different positions of the point source, giving the bottom-left part of $A_{12}$ (also see Fig. 3a and c). It can be seen that the magnitude of the elements decreases rapidly with increasing distance from the boundary. Therefore, we can truncate the off-diagonal blocks and keep only $t\ll N$ points, so that $A_{12}$ can be implemented as a $t\times t$ matrix-vector multiplication using only the first or last $t$ points of each edge of the subdomain.
Due to the symmetry of the angular spectrum kernel, the off-diagonal elements of $V_{11}$ ($V_{22}$) are equal to $(A_{12}+A_{12}^{T})$ (see Fig. 3d). In other words, we can use the same operator (up to a minus sign) to compute both the transfer between subdomains and the wrapping correction inside each subdomain.
Note that this approach also eliminates the wrapping artefacts present in the original system (the corners in Fig. 2d). We therefore eliminate the need for expensive zero padding or perfectly matched layers [41, 10], and instead use a thin anti-reflective boundary [3], reducing the computational costs and memory requirements.
@@
Figure 5: Using truncated blocks in $V$ to reconstruct AA. (a) Matrix representation of AA decomposed over two subdomains (demarcated by the dashed grey lines), but with truncated blocks. (b) $L$ decomposed over the subdomains remains unchanged and is the same as in Fig.3b. (c) $V$ with truncated blocks for communication (off-diagonal) and removing wrapping artefacts (off-diagonal elements of the diagonal blocks), visible as square blocks of side $t=6$ in the corners of subdomain blocks. Note that even though the colour scale is logarithmic in both the positive and negative directions from 0, the blocks are hardly visible, showing that the effect of the truncation is small.
@@
Figure 5 is a reproduction of Fig. 3 but with truncated blocks for communication between subdomains (the off-diagonal blocks $A_{12}$ and $A_{21}$) and for removing the wrapping artefacts (the off-diagonal elements of $V_{11}$ and $V_{22}$). The operator $A$ obtained using these truncated blocks is shown in Fig. 5a. The decomposed operator $L$ (Eq. (8)) shown in Fig. 5b remains unchanged irrespective of the truncation and is the same as in Fig. 3b. Figure 5c shows $V$ obtained using the truncation instead of the full blocks. The effect of the truncation is hardly visible in Fig. 5c as square blocks of size $t=6$, indicating that the effect of the truncation is small. In Section 4.2.1, we analyse quantitatively the effect of the truncation parameter $t$ on the accuracy of the simulation.